OpenStudy (anonymous):

A woman can walk 2mph faster down a trail than she can returning up the trail. It takes her 2 hr to get down and 4 hr to get back up. What is her speed down the trail?

OpenStudy (jdoe0001):

so she goes up the trail and then down the trail now the trail doesn't get longer or shorter, thus the distance on both cases is the same so we can say that \(\begin{array}{llll} &\bf distance& \bf rate &\bf time\\ \textit{up the trail}&d& r& 4\\ \textit{down the trail}&d& r+2& 2 \end{array}\) distance = rate x time so UP distance and DOWN distance are the same, or equal to each other, thus d = d rt = rt => r*4 = (r+2) * 2 # solve for "r" to find the the rate when she's going UP whatever that gives you, the DOWN rate is THAT MUCH + 2